OpenStudy (aimmebgood):

Factor completely x^(2) + 2xy + y^(2) - a^(2).

OpenStudy (anonymous):

@aimmeBgood i have no idea how to solve this so i'm here to learn!! And @satellite73 most likely understands this so i'm here to understand too! :)

OpenStudy (anonymous):

\[x^2+2xy+y^2\] is a perfect square it is \((x+y)^2\)

OpenStudy (anonymous):

ohhh you take it apart? into pieces? and factor what you can?

OpenStudy (anonymous):

anytime you see something that looks like \(a^2+2ab+b^2\) you have a perfect square so now you can start with \[(x+y)^2-a^2\] which is itself the difference of two squares

OpenStudy (anonymous):

oh and i see how you got (x+y)^2 for the first three terms... :) so would the answer be 

OpenStudy (anonymous):

you can go further

OpenStudy (anonymous):

ohhh :) how?

OpenStudy (anonymous):

because now you have the difference of two squares \[A^2-B^2=(A+B)(A-B)\]

OpenStudy (aimmebgood):

but where does the 2 go? i dont understand.

OpenStudy (anonymous):

replace \(A\) by \((x+y)\) and \(B\) by \(a\)

OpenStudy (anonymous):

the 2 does not "go" anywhere, it is hidden in \((x+y)^2\)

OpenStudy (anonymous):

if you multiply out you get \((x+y)^2=(x+y)(x+y)=x^2+xy+yx+y^2=x^2+2xy+y^2\)

OpenStudy (anonymous):

your final answer should look like \((x+y+a)(x+y-a)\)

OpenStudy (aimmebgood):

that makes sense. i get it now. thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh wow! that's cool :) that makes sense!! :) @satellite73 strikes again! :) And thanks for the explanation! (even tho it wasn't my question, i understand it now!) :D